# Balancing Confidentiality and Transparency for Blockchain-based Process-Aware Information Systems

### v2

@@ 5.3 Execution Analysis
Table 2 shows the execution time and costs registered during the execution of the X-ray diagnostic process. For the time evaluation, we ran a separate process instance for each path in the control flow (each corresponding to a different decision in correspondence with the exclusive-choice gateways) five times (thus totalling twenty executions). The registered values are reported in Table 2(a). We employed the Sepolia testnet as the Blockchain Platform, and passed input data based on real-world files in the healthcare domain upon anonymization. The dataset and the Sepolia transaction handles are openly accessible at github.com/apwbs/CONFETTY.
[…]
| min | max | avg | |
| Configure | |||
| Instantiate | |||
| Transact | |||
| Inspect | |||
| Functionality | Used gas units |
| Instantiate | |
| Transact | |
| Inspect | |
| Deployment | |
| Total |
@@
The second part of our experimental evaluation illustrates the time and cost needed by CONFETTY to run three different processes from the literature, namely the aforementioned X-ray [9], retail [7] and incident management [10]. Figure 9 illustrates the results, averaging overruns that took different branches in correspondence with exclusive choice gateways. We gathered data using a local RPC as a stub for the Programmable Blockchain Platform. The three processes exhibit similar trends from both standpoints. The Incident management process is the heaviest and most expensive one, yet exhibiting an increment of for the cumulative time and for the cumulative cost with respect to the lower bound of the X-ray process. This increment depends on the number of process participants ( in the Incident management against the of the X-ray process), thus leading to the execution of more steps.
@@ 6 Conclusions
In this work, we presented CONFETTY, a blockchain-based architecture for PAISs. CONFETTY allows for transparency and public enforcement of process execution while preserving confidentiality of sensitive data. A public state saved on chain maintains process data used for enforcement and auditability purposes. The evolution of the public state is mediated by smart contract implementing the business process control- and data-flow logic. We use ABE encryption to hide sensitive information and grant access to it only to authorized parties based on their roles in the process, encoded as attributes in access policies and embedded directly in their personal decryption keys. To evaluate the CONFETTY architecture we have analyzed its threat model and measured its time and cost performance through the execution of an X-ray examination choreography process. Future work includes extensive performance and scalability analyses against synthetic and real-world benchmarks. We also envision a field study to assess the practical adoption and usability with real-world stakeholders. Another interesting avenue for further investigations is the full decentralization of the architecture, to remove the existing points of trust residing in off-chain components. Finally, we foresee a study and design of a solution to enable automated decision support via computation held on encrypted data.

### v3

@@ 5.3 Execution Analysis
5.3.1 Execution Performance and Expenditure.
We gauged the time and cost of our prototypical implementation against data simulating the execution and message exchange of processes from the literature. Figure 7 illustrates the results, averaging over runs that took the longest path. We gathered data using a local RPC as a stub for the Programmable Blockchain Platform.
@@
Let us now compare the above results with two other processes from the literature, namely retail [6] and incident management [8]. The X-ray, incident and retail processes have similar structures in the number of messages (10, 13 and 12, respectively), gateways (5, 3 and 2) and involved participants (4, 5 and 3) and exhibit similar trends from both runtime and cost standpoints. The incident management process is the slowest and most expensive one with an increase in cumulative time and cost compared to the X-ray process, respectively. The experiments show that CONFETTY performs comparably to literature tools and scenarios. Indeed, ChorChain [8] required approximately million GU for the incident management process, while MultiChain [6] required around million GU for the retail process. Notice that such approaches only focus on public execution and enforcement without incorporating any confidential mechanisms and that they deploy a contract for each new instance.
5.3.2 Scalability Assessment.
Lastly, we analyzed the scalability of CONFETTY. To this end, we considered the dependence of time and cost taking as independent variables the dimensions of the process affecting the input size for our tool. In this section, we focus on the number of process participants and the choreography model size due to space restrictions as they evidence the most salient results. Further experiments including other variables like the number of gateways (parallel and exclusive) and the message payload size can be found in the online repository. The default values for the variables adhere to the example provided in Fig. 1. For every test, we let one variable change the value and keep the others fixed, assigned with the respective defaults. Figures 8 and 9 dissect the registered timings and costs to serve the different functionalities concurring to the total by stacking the respective subtotals.
[…]
displays the results observed varying the process model size. We progressively increased it by sequentially concatenating the whole choreography diagram multiple times to derive a new one, thus scaling the size up to ten times the original one. Figure 9(a) shows a linear growth in the runtime of the transact and inspect functionalities, without a significant impact on the remaining one, instantiate. The execution time increases with the number of messages due to additional read and write operations. Figure 9(b) shows the cost increase, with a linear growth registered for the instantiate and transact functionalities. More elements increase on-chain storage needs and hence gas cost. Read operations remain constant as they only handle one-time key management on-chain (hence the flat cost associated with inspect).
@@ 6 Conclusions
In this work, we presented CONFETTY, a blockchain-based architecture for PAISs. CONFETTY enables transparency and public enforcement for process execution while preserving the confidentiality of sensitive exchanged data. A public state saved on-chain maintains process data for enforcement and auditability purposes. The evolution of the public state is mediated by smart contracts implementing the business process control- and data-flow logic. We use MA-ABE encryption to hide sensitive information and grant access only to authorized parties based on their roles, encoded as attributes in access policies and embedded directly in their decryption keys. We confirmed its security guarantees with a threat model analysis, and demonstrated the time and cost performance of our implemented prototype with processes taken from the literature.
In future work, we envision a field study to assess the practical adoption and usability with real-world stakeholders. Another interesting avenue is the full decentralization of the architecture (to remove the existing points of trust residing in off-chain components) and governance (to cater for multiple process owners during the instantiation phase). Conducting a security analysis of our approach against stronger adversarial models is an intriguing future endeavor, too. Finally, we envision a study and design of a solution to enable automated decision support via computation held on encrypted data.
6.0.1 Acknowledgements
This work was partly funded by projects PINPOINT (B87G22000450001) under the PRIN MUR program, and Health-e-Data, funded by the EU-NGEU under the Cyber 4.0 NRRP MIMIT programme.